# The N3LO Twist-2 Matching of Helicity TMDs and SIDIS $q_{\ast}$ Spectrum

### v3

@@ 4.4.2 Time-like results in $\overline{\text{MS}}$}
$\Delta P_{qq}^{T,\text{s}(2)}=C_{F}^{3}\bigg[\frac{1}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}-56H_{0,1,0}+96H_{1,0,0}-208H_{0,0,0,1}-64H_{0,0,1,1}-48H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z}{z-1}\left(-192\zeta_{2}H_{1}-800H_{1}+64H_{0,1,0}\right)+\frac{z}{z-1}\left(-144\zeta_{2}H_{0}-72H_{0}-12\zeta_{2}-192\zeta_{3}-324\zeta_{4}-352\zeta_{2}H_{0,-1}+144\zeta_{2}H_{0,0}+176H_{0,0}-96H_{0,0,0}-64H_{0,-1,-1,0}+128H_{0,-1,0,0}-96H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(-16\zeta_{2}H_{0}+128\zeta_{3}H_{0}+10H_{0}+204\zeta_{2}+48\zeta_{3}+252\zeta_{4}+160\zeta_{2}H_{0,-1}+208\zeta_{2}H_{0,0}+172H_{0,0}+528H_{0,0,0}-320H_{0,-1,-1,0}+192H_{0,-1,0,0}+288H_{0,0,-1,0}-592H_{0,0,0,0}-192H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}+88H_{0,1,0}+96H_{1,0,0}-240H_{0,0,0,1}-64H_{0,0,1,1}-144H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z^{3}}{z-1}\left(432\zeta_{2}H_{0}+114H_{0}+12\zeta_{2}+336\zeta_{3}-28\zeta_{4}-352\zeta_{2}H_{0,-1}+48\zeta_{2}H_{0,0}-228H_{0,0}+240H_{0,0,0}-64H_{0,-1,-1,0}-160H_{0,0,-1,0}-112H_{0,0,0,0}-128H_{0,0,1,0}+\frac{1}{z+1}+302\right)+\frac{1}{z+1}\left(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-12H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-64H_{0,-1,0}+64H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\right)+\frac{1}{(z-1)(z+1)}\left(16\zeta_{2}H_{0}-64\zeta_{3}H_{0}-64H_{0}-204\zeta_{2}-192\zeta_{3}-44\zeta_{4}+160\zeta_{2}H_{0,-1}+112\zeta_{2}H_{0,0}-224H_{0,0}-96H_{0,0,0}-320H_{0,-1,-1,0}+64H_{0,-1,0,0}+96H_{0,0,-1,0}-176H_{0,0,0,0}-128H_{0,0,1,0}+302\right)+\frac{z}{z+1}\left(-576H_{-1}\zeta_{2}-192H_{-1,0}-312H_{0,1}-64H_{-1,0,0}+768H_{-1,0,1}-320H_{0,-1,0}-96H_{0,0,1}\right)+\frac{z^{2}}{z+1}\left(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-300H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-160H_{0,-1,0}-256H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\right)+144H_{1,0}+192H_{-1,-1,0}+z\left(192H_{-1,-1,0}-144H_{1,0}\right)\bigg]+C_{F}^{2}N_{f}\bigg[\frac{668}{3}H_{1}+\frac{1}{z-1}\left(\left(-\frac{2104}{9}\right)H_{0,1}+\left(-\frac{464}{3}\right)H_{0,1,0}+\left(-\frac{580}{9}\right)H_{1,0}-160\zeta_{2}H_{0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{5048}{9}\right)H_{0,0}+\left(-\frac{496}{3}\right)H_{0,0,1}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{4772}{9}\zeta_{2}+\frac{7868}{9}H_{0}+\left(-\frac{112\zeta_{2}}{3}\right)H_{0}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-152H_{0,0,0}+\frac{1}{z+1}-\frac{1727}{3}\right)+\frac{z}{z-1}\left(\frac{220}{3}H_{0,1}+\frac{280}{3}H_{1,0}+144H_{0,1,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{580}{9}\right)H_{1,0}+\left(-\frac{32}{3}\right)H_{0,1,0}+\frac{1124}{9}H_{0,1}+160\zeta_{2}H_{0,1}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{80}{3}\right)\zeta_{3}+\frac{20}{9}H_{0}+\frac{56}{3}H_{0,0,0}+\frac{760}{9}H_{0,0}+\frac{1672}{9}\zeta_{2}-64H_{0}\zeta_{2}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+64H_{0,0,1}+\frac{1}{z+1}-539\right)$ (128)
$\frac{z^{3}}{z-1}\left(\left(-\frac{7664}{9}\right)H_{0}+\left(-\frac{5092}{9}\right)\zeta_{2}+\frac{112}{3}\zeta_{3}+\frac{368}{3}H_{0,0,1}+\frac{1624}{3}H_{0,0}+\frac{176\zeta_{2}}{3}H_{0}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+120H_{0,0,0}+\frac{1}{z+1}+\frac{1727}{3}\right)+\frac{1}{z+1}\left(\left(-\frac{512}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,0,1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{-1}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{1576}{9}\right)H_{0,0}+\left(-\frac{1352}{9}\right)\zeta_{2}+\left(-\frac{280}{3}\right)H_{0,0,0}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{184}{9}H_{0}+64H_{0}\zeta_{2}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-64H_{0,0,1}+539\right)+\left(-\frac{128z}{3(z+1)}\right)H_{-1,0}+\frac{z^{2}}{z+1}\left(\left(-\frac{512}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,0,1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{-1}\right)+400H_{1}\zeta_{2}-52H_{1,1}+48H_{0,1,1}-440H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-64H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-176H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}+z\left(\left(-\frac{668}{3}\right)H_{1}-400H_{1}\zeta_{2}+52H_{1,1}-32H_{0,1,1}+440H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}-64H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-176H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}\right)-32H_{0,1,1,1}\bigg]+C_{A}^{2}C_{F}\bigg[\frac{1}{z-1}\left(\frac{592}{3}H_{1}+32H_{1}\zeta_{2}-144H_{1}\zeta_{3}+16\zeta_{2}H_{1,0}+76H_{1,0,0}-80H_{0,-1,-1,0}+36H_{0,1,0,0}-96H_{1,0,-1,0}+48H_{1,0,0,0}-64H_{1,0,0,1}+64H_{1,1,0,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1184}{3}\right)H_{1}-64H_{1}\zeta_{2}-64H_{1,0,0}+32H_{0,-1,-1,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1744}{9}\right)\zeta_{2}+\left(-\frac{722}{9}\right)H_{0,0}+\left(-\frac{1258}{27}\right)H_{0}+\frac{160}{3}\zeta_{3}+\frac{344}{3}H_{0,-1,0}-100H_{0}\zeta_{2}+7\zeta_{4}+60\zeta_{2}H_{0,0}+100H_{0,0,1}+72H_{0,-1,0,0}+96H_{0,-1,0,1}-8H_{0,0,-1,0}-60H_{0,0,0,1}+\frac{1}{z+1}+\frac{5333}{27}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{416}{3}\right)H_{0,0,0}+\left(-\frac{1378}{27}\right)H_{0}+\left(-\frac{4}{3}\right)H_{0,0,1}+\frac{328}{3}H_{0,-1,0}+\frac{1504}{9}\zeta_{2}+\frac{332\zeta_{2}}{3}H_{0}-144H_{0}\zeta_{3}+112\zeta_{3}-121\zeta_{4}-68\zeta_{2}H_{0,0}-130H_{0,0}-56H_{0,-1,0,0}-160H_{0,-1,0,1}-40H_{0,0,-1,0}+40H_{0,0,0,0}+28H_{0,0,0,1}+\frac{1}{z+1}+\frac{923}{27}\right)+\frac{z^{2}}{z-1}\left(\frac{592}{3}H_{1}+32H_{1}\zeta_{2}-144H_{1}\zeta_{3}+16\zeta_{2}H_{1,0}+76H_{1,0,0}-80H_{0,-1,-1,0}+28H_{0,1,0,0}-96H_{1,0,-1,0}+48H_{1,0,0,0}-64H_{1,0,0,1}+64H_{1,1,0,0}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{376}{3}\right)H_{0,-1,0}+\left(-\frac{2266}{27}\right)H_{0}+\frac{238}{9}H_{0,0}+\frac{224}{3}\zeta_{2}+188H_{0}\zeta_{2}+48\zeta_{3}-31\zeta_{4}+36\zeta_{2}H_{0,0}-188H_{0,0,1}+88H_{0,-1,0,0}+96H_{0,-1,0,1}+8H_{0,0,-1,0}-36H_{0,0,0,1}+\frac{1}{z+1}-\frac{5333}{27}\right)+\frac{1}{z+1}\left(\left(-\frac{1312}{9}\right)H_{-1,0}+\left(-\frac{224}{3}\right)H_{0,1}+\frac{56}{3}H_{-1,0,0}+\frac{656}{3}H_{-1,0,1}+\left(-\frac{560\zeta_{2}}{3}\right)H_{-1}-192H_{-1}\zeta_{3}+256\zeta_{2}H_{-1,-1}-176\zeta_{2}H_{-1,0}-120\zeta_{2}H_{0,-1}-40\zeta_{2}H_{0,1}-128H_{-1,-1,0,0}-256H_{-1,-1,0,1}+48H_{-1,0,0,0}+128H_{-1,0,0,1}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{560}{3}\right)\zeta_{3}+\left(-\frac{3442}{27}\right)H_{0}+\left(-\frac{80}{3}\right)H_{0,0,0}+\left(-\frac{8}{3}\right)H_{0,-1,0}+\frac{92}{3}H_{0,0,1}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-48\zeta_{2}-112H_{0}\zeta_{3}-63\zeta_{4}-28\zeta_{2}H_{0,0}-162H_{0,0}-40H_{0,-1,0,0}-160H_{0,-1,0,1}-24H_{0,0,-1,0}+24H_{0,0,0,0}+4H_{0,0,0,1}-\frac{923}{27}\right)+\frac{z}{z+1}\left(\left(-\frac{448}{3}\right)H_{0,1}+\left(-\frac{160}{3}\right)H_{-1,0}-256H_{-1}\zeta_{2}-16\zeta_{2}H_{0,1}+96H_{-1,0,0}+320H_{-1,0,1}\right)$ (129)
[…]

### v4

@@ 4.4.2 Time-like results in $\overline{\text{MS}}$}
$\Delta P_{qq}^{T,\text{s}(2)}=C_{F}N_{f}^{2}\bigg[\left(-\frac{416}{9}\right)H_{1}+\left(-\frac{152}{9}\right)H_{0,1}+\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{16}{3}H_{0,0,1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,1}+\frac{152}{9}\zeta_{2}+\frac{64}{3}\zeta_{3}+\frac{1}{z-1}\left(\frac{8}{3}H_{0,0}+\frac{112}{27}H_{0}+\frac{16\zeta_{2}}{3}H_{0}\right)+\frac{z}{z-1}\left(\left(-\frac{64}{3}\right)H_{0,0}+\frac{320}{9}H_{0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{128}{3}\right)H_{0}+\frac{152}{9}H_{0,0}+\left(-\frac{16\zeta_{2}}{3}\right)H_{0}\right)+40H_{1,0}+16H_{0,1,0}+z\left(\left(-\frac{296}{9}\right)H_{0,1}+\left(-\frac{40}{3}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{16}{3}H_{0,0,1}+\frac{16}{3}H_{0,1,1}+\frac{64}{3}\zeta_{3}+\frac{296}{9}\zeta_{2}+\frac{416}{9}H_{1}-40H_{1,0}+16H_{0,1,0}+\frac{1952}{27}\right)-\frac{1936}{27}\bigg]+\left[\frac{1}{1-z}\right]_{+}\bigg[C_{F}\left(\left(-\frac{1072}{9}\right)\zeta_{2}+\frac{88}{3}\zeta_{3}+88\zeta_{4}+\frac{490}{3}\right)C_{A}^{2}+\left(\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}-\frac{836}{27}\right)C_{F}N_{f}C_{A}-\frac{16}{27}C_{F}N_{f}^{2}+C_{F}^{2}N_{f}\left(32\zeta_{3}-\frac{110}{3}\right)\bigg]+\delta(1-z)\bigg[\left(-32\zeta_{3}\zeta_{2}+18\zeta_{2}+68\zeta_{3}+144\zeta_{4}-240\zeta_{5}+\frac{29}{2}\right)C_{F}^{3}+\left(\left(-\frac{136}{3}\right)\zeta_{3}+\frac{20}{3}\zeta_{2}+\frac{116}{3}\zeta_{4}-23\right)N_{f}C_{F}^{2}+\left(\left(-\frac{16}{9}\right)\zeta_{3}+\frac{80}{27}\zeta_{2}-\frac{17}{9}\right)N_{f}^{2}C_{F}+C_{A}^{2}\left(\left(-\frac{1552}{9}\right)\zeta_{3}+\frac{4496}{27}\zeta_{2}-5\zeta_{4}+40\zeta_{5}-\frac{1657}{36}\right)C_{F}+C_{A}\left(\left(\left(-\frac{494}{3}\right)\zeta_{4}+\left(-\frac{410}{3}\right)\zeta_{2}+\frac{844}{3}\zeta_{3}+16\zeta_{2}\zeta_{3}+120\zeta_{5}+\frac{151}{4}\right)C_{F}^{2}+N_{f}\left(\left(-\frac{1336}{27}\right)\zeta_{2}+\frac{200}{9}\zeta_{3}+2\zeta_{4}+20\right)C_{F}\right)\bigg]$
$C_{F}^{3}\left[\frac{1}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}-56H_{0,1,0}+96H_{1,0,0}-208H_{0,0,0,1}-64H_{0,0,1,1}-48H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z}{z-1}\left(-192\zeta_{2}H_{1}-800H_{1}+64H_{0,1,0}\right)+\frac{z}{z-1}\left(-144\zeta_{2}H_{0}-72H_{0}-12\zeta_{2}-192\zeta_{3}-324\zeta_{4}-352\zeta_{2}H_{0,-1}+144\zeta_{2}H_{0,0}+176H_{0,0}-96H_{0,0,0}-64H_{0,-1,-1,0}+128H_{0,-1,0,0}-96H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(-16\zeta_{2}H_{0}+128\zeta_{3}H_{0}+10H_{0}+204\zeta_{2}+48\zeta_{3}+252\zeta_{4}+160\zeta_{2}H_{0,-1}+208\zeta_{2}H_{0,0}+172H_{0,0}+528H_{0,0,0}-320H_{0,-1,-1,0}+192H_{0,-1,0,0}+288H_{0,0,-1,0}-592H_{0,0,0,0}-192H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}+88H_{0,1,0}+96H_{1,0,0}-240H_{0,0,0,1}-64H_{0,0,1,1}-144H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z^{3}}{z-1}\left(432\zeta_{2}H_{0}+114H_{0}+12\zeta_{2}+336\zeta_{3}-28\zeta_{4}-352\zeta_{2}H_{0,-1}+48\zeta_{2}H_{0,0}-228H_{0,0}+240H_{0,0,0}-64H_{0,-1,-1,0}-160H_{0,0,-1,0}-112H_{0,0,0,0}-128H_{0,0,1,0}+\frac{1}{z+1}+302\right)+\frac{1}{z+1}\left(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-12H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-64H_{0,-1,0}+64H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\right)+\frac{1}{(z-1)(z+1)}\bigg(16\zeta_{2}H_{0}-64\zeta_{3}H_{0}-64H_{0}-204\zeta_{2}-192\zeta_{3}-44\zeta_{4}+160\zeta_{2}H_{0,-1}+112\zeta_{2}H_{0,0}-224H_{0,0}-96H_{0,0,0}-320H_{0,-1,-1,0}+64H_{0,-1,0,0}+96H_{0,0,-1,0}-176H_{0,0,0,0}-128H_{0,0,1,0}+302\bigg)+\frac{z}{z+1}\left(-576H_{-1}\zeta_{2}-192H_{-1,0}-312H_{0,1}-64H_{-1,0,0}+768H_{-1,0,1}-320H_{0,-1,0}-96H_{0,0,1}\right)+\frac{z^{2}}{z+1}\bigg(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-300H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-160H_{0,-1,0}-256H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\bigg)+144H_{1,0}+192H_{-1,-1,0}+z\left(192H_{-1,-1,0}-144H_{1,0}\right)\right]+C_{F}^{2}N_{f}\bigg[\frac{668}{3}H_{1}+\frac{1}{z-1}\left(\left(-\frac{2104}{9}\right)H_{0,1}+\left(-\frac{464}{3}\right)H_{0,1,0}+\left(-\frac{580}{9}\right)H_{1,0}-160\zeta_{2}H_{0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{5048}{9}\right)H_{0,0}+\left(-\frac{496}{3}\right)H_{0,0,1}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{4772}{9}\zeta_{2}+\frac{7868}{9}H_{0}+\left(-\frac{112\zeta_{2}}{3}\right)H_{0}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-152H_{0,0,0}+\frac{1}{z+1}-\frac{1727}{3}\right)+\frac{z}{z-1}\left(\frac{220}{3}H_{0,1}+\frac{280}{3}H_{1,0}+144H_{0,1,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{580}{9}\right)H_{1,0}+\left(-\frac{32}{3}\right)H_{0,1,0}+\frac{1124}{9}H_{0,1}+160\zeta_{2}H_{0,1}\right)+\frac{z^{2}}{z-1}\bigg(\left(-\frac{80}{3}\right)\zeta_{3}+\frac{20}{9}H_{0}+\frac{56}{3}H_{0,0,0}+\frac{760}{9}H_{0,0}+\frac{1672}{9}\zeta_{2}-64H_{0}\zeta_{2}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+64H_{0,0,1}+\frac{1}{z+1}-539\bigg)+\frac{z^{3}}{z-1}\bigg(\left(-\frac{7664}{9}\right)H_{0}+\left(-\frac{5092}{9}\right)\zeta_{2}+\frac{112}{3}\zeta_{3}+\frac{368}{3}H_{0,0,1}+\frac{1624}{3}H_{0,0}+\frac{176\zeta_{2}}{3}H_{0}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+120H_{0,0,0}+\frac{1}{z+1}+\frac{1727}{3}\bigg)$
[…]